# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ 1.3 Proof strategy
In this subsection, we discuss the high-level intuition of the strategy we will follow, building from the case of $d=1$. Throughout we condition on the survival event. We first assume the jump is spherically symmetric. For one-dimensional BRW, we expect the following parity relation between maximum and FPT. Recall that the asymptotic for the maximum $M_{n}$ at time $n$ is
@@
$\tau_{x}=m_{1}(x)+O_{\mathbb{P}}(1),\quad\text{ where }\quad m_{1}(x):=\frac{x}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log x,$ (9)
which we will prove in Proposition 7 below in Section 2.1. Recall (3). Let us decompose $A(x)=m_{1}(x)+m_{2}(x)$, where
$m_{2}(x):=A(x)-m_{1}(x)=\frac{d-1}{2c_{2}c_{1}}\,\log x.$ (10)
For a general dimension $d\geqslant 1$, we shall explain the extra term $m_{2}(x)$.
@@
A key step in these bounds is Proposition 8, which keeps track of the number of particles that are near the frontier (say, a distance of $\log n$ away) for a one-dimensional BRW. Roughly speaking, there are around $xe^{c_{2}x}$ many particles that are of a distance $x$ behind the frontier, where $x\ll\log n$. This amounts to generalizing the proof of the asymptotic of the maximum (8); see [13]. A notable distinction is that unfortunately, one cannot conclude from the second moment method the lower bound for the number of particles near the frontier with overwhelming probability. Instead, one only concludes the lower bound with a uniformly positive probability. This gap can be filled using Theorem 2, which shows that a uniformly positive probability would already be sufficient.
A particularly nice feature of the spherical symmetry of the jumps is that the displacements in different directions are approximately independent (likewise, the same result Theorem 1 holds if we instead assume independent centered jumps in the $d$ directions). In particular, at time $n$, one roughly expects a proportion $O(n^{-(d-1)/2})$ of the particles that lie within unit distance from the origin in the last $d-1$ dimensions. In other words, one expects that the first passage event is realized around the time when $x^{(d-1)/2}$ particles reach a distance of $x$ for the $\mathbb{R}^{d}$-valued BRW projected onto the first dimension; see Figure 2 below. On the other hand, by (8), the extra time of $m_{2}(x)$ from (10) pushes the frontier ahead by a distance of approximately $c_{1}m_{2}(x)=(d-1)(\log x)/(2c_{2})$, which amounts to having approximately $x^{(d-1)/2}$ extra particles in view of Proposition 8.
@@
If the jump distribution is not spherically symmetric, applying similar arguments above (with Proposition 8 replaced by Theorem B of [11] and Proposition 10 by Proposition 17; we omit the details here) will lead to
@@
where the corresponding upper and lower bounds for $\tau_{x}$ are separated by an order of $\log x$. To pin down the coefficient of the logarithm correction term, we extend the arguments for the maximum of one-dimensional BRW ([13]) with a modified barrier. The modified barrier event lies in $\mathbb{R}^{d}$ and hence results on random walks in cones ([23]) are applicable. Note that the same proof also applies for the spherically symmetric case, but would yield less precise asymptotics compared to (4) (an $o_{\mathbb{P}}(\log x)$ instead of an $O_{\mathbb{P}}(\log\log x)$).
@@ 2 Proofs of main results
This section is devoted to the proofs of the main results. We will first analyze the simple case of $d=1$ in Section 2.1, where a precise asymptotic of the FPT follows directly from inverting the maxima asymptotic (8). We then collect the necessary tools in Sections 2.2.1 and 2.2.2, and prove Theorems 1 and 2 in Sections 2.2.3 and 2.3 respectively. The case of non-spherically symmetric jumps (Theorem 3) will be addressed in Section 2.4, where we also summarize preliminary results on random walks that are constrained to stay in a cone. Section 2.5 presents the proofs of Corollaries 4 and 5.
[…]
@@ 2.1 Warming up in dimension one
For $\beta>0$ and $n\in\mathbb{N}$ we define the event
$G_{n,\beta}:=\bigcup_{v\in V_{n}}\bigcup_{0\leqslant k\leqslant n}\left\{\eta_{v,n}(k)\geqslant\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\right\}.$
Lemma 6 (Lemma 2.4 of [13]).
Assume conditions (A1), (A2), and (A4), and assume that the jump distribution is non-lattice. It holds that $\mathbb{P}(G_{n,\beta})\ll\beta e^{-c_{2}\beta}$.
@@
Suppose that a one-dimensional BRW satisfies conditions (A1), (A2), and (A4), and that the jump distribution is non-lattice. We have
$\tau_{x}=m_{1}(x)+O_{\mathbb{P}}(1)=\frac{x}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log x+O_{\mathbb{P}}(1).$
@@
For the lower bound, we fix $\varepsilon>0$ and pick $\beta$ large enough such that $\mathbb{P}(G_{n,\beta})<\varepsilon$, by means of Lemma 6. In other words, with a probability larger than $1-\varepsilon$, all trajectories of the BRW stay below the curve
@@
$\max_{1\leqslant k\leqslant\tau_{x}-C_{1}}\Big{(}\frac{km_{\tau_{x}}}{\tau_{x}}+\beta+\frac{4}{c_{2}}(\log(\min(k,\tau_{x}-k)))_{+}\Big{)}\leqslant x.$
This yields $\mathbb{P}(\tau_{x}\leqslant m_{1}(x)-C_{1})\leqslant\mathbb{P}(G_{n,\beta})<\varepsilon$, proving the lower bound.
For the upper bound, we fix $\varepsilon>0$. Pick $C_{2}>0$ such that $\mathbb{P}(M_{m_{1}(x)+C_{2}}>x)>1-\varepsilon/3$ by (8). Using the same analysis for the lower bound, there exists $C_{3}>0$ such that $\mathbb{P}(M_{m_{1}(x)+C_{2}}>x+C_{3})<\varepsilon/3$. Therefore, with probability $>1-2\varepsilon/3$, there is $v\in V_{m_{1}(x)+C_{2}}$ such that $x\leqslant\eta_{v,m_{1}(x)+C_{2}}(m_{1}(x)+C_{2})\leqslant x+C_{3}$. Since the distribution of $\xi$ is non-lattice, by the Theorem 2 of [42], the random walk initiated from $v$ is recurrent, and hence there exists $C_{4}>0$ such that the descendants of $v$ reaches $[x-1/2,x+1/2]$ with probability $>1-\varepsilon/3$ within time $C_{4}$. Combining the above leads to $\mathbb{P}(\tau_{x}>m_{1}(x)+C_{2}+C_{4})<\varepsilon$, as desired. ∎
@@ 2.2 Proof of Theorem 1
We now present the proof of Theorem 1 for $d\geqslant 2$, conditioned on Theorem 2. The proof of Theorem 2 will be the focus of the next section. The following auxiliary results are necessary as explained in our proof strategy.
@@ 2.2.1 The number of particles near maximum
The goal of this section is to prove the following result on the number of particles near the frontier of the BRW, which might be of independent interest.
@@
Assume (A1)–(A4). There exists $L>0$ depending only on the law of the BRW such that the following holds conditioned upon survival. Given any $\varepsilon>0$, there exists $C>0$ independent from $n$ and $x$ such that uniformly for $n$ large enough and for $x\in[2,\sqrt{n}]$,
(#$\mathbb{P}\left(\#\{v\in V_{n}:\eta_{v,n}(n)\geqslant m_{n}-x\}>Cxe^{c_{2}x}\mid S\right)<\varepsilon$ (12)
@@
(#$\mathbb{P}\left(\#\{v\in V_{n}:\eta_{v,n}(n)\geqslant m_{n}-x\}>\frac{1}{C}xe^{c_{2}x}\mid S\right)>\frac{1}{L}.$ (13)
@@
Proposition 8 may be compared against [11, 45], where it is proven that for a location that is $\Omega(n)$ away from $m_{n}$, the number of particles therein has the same order as its expectation. Proposition 8 shows that this is not the case for locations that are $O(\sqrt{n})$ near $m_{n}$. On the other hand, the number of particles that are $O(\sqrt{n})$ away from the origin can be described by the BRW central limit theorem [10]. An interesting question would be to analyze the phase transitions between those regimes. We also expect that (13) holds if we replace its right-hand side by $1-\varepsilon$, but the current version suffices for our purpose.
Following the setup of Section 2.2 of [13], we introduce $\lambda_{n}$ which is the value of $\lambda$ where the supremum of (1) is attained with $x=m_{n}/n$. By (15) of [13], it holds $0\leqslant c_{2}-\lambda_{n}\ll(\log n)/n$. Let $\mathbb{Q}^{(n)}$ be defined by
$\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}^{(n)}}:=e^{-\lambda_{n}(\eta_{v,n}(n)-m_{n})-nI(m_{n}/n)}\asymp n^{3/2}\rho^{-n}e^{-\lambda_{n}(\eta_{v,n}(n)-m_{n})}.$ (14)
It follows that under $\mathbb{Q}^{(n)}$, $\{\eta_{v,n}(k)-km_{n}/n\}_{k=0,\dots,n}$ is a mean zero random walk. Let
$Q_{n,\beta}:=\left\{v\in V_{n}:\text{ for any }0\leqslant k\leqslant n,\leavevmode\nobreak\ \eta_{v,n}(k)<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\right\}$
@@
$g(n,\beta,x):={\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta_{v,n}(n)\geqslant m_{n}-x\}]}.$ (15)
@@
$\mathbb{E}\left[\#\{v\in V_{n}:\eta_{v,n}(n)\geqslant m_{n}-x\}{\mathbbm{1}}_{G_{n,\beta}^{c}}\right]\ll\beta(x+\beta)e^{c_{2}x}.$
@@
$\mathbb{P}\left(G_{n,\beta}^{c}\cap\left\{\#\{v\in V_{n}:\eta_{v,n}(n)\geqslant m_{n}-x\}>L\beta(x+\beta)e^{c_{2}x}\right\}\right)<\frac{\varepsilon}{2}.$
@@
Consider a fixed $v\in V_{n}$. We have
$\hskip 14.22636pt\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta_{v,n}(n)\geqslant m_{n}-x\}]$
$=\rho^{n}\mathbb{P}(v\in Q_{n,\beta},\leavevmode\nobreak\ \eta_{v,n}(n)\geqslant m_{n}-x)$
$=\sum_{j=0}^{x+\beta-1}\rho^{n}\mathbb{P}(v\in Q_{n,\beta},\leavevmode\nobreak\ \eta_{v,n}(n)\in[m_{n}+\beta-j-1,m_{n}+\beta-j)).$
By substituting $i=\beta-j$, the latter probability is precisely $\chi_{n,n}^{\mathbb{P}}(i)$ defined below (15) in [13]. It is proved in (18) therein the estimate
@@
$\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta_{v,n}(n)\geqslant m_{n}-x\}]\ll\sum_{j=0}^{x+\beta}\beta(j+2)e^{-c_{2}(\beta-j)}\ll\beta(x+\beta)e^{c_{2}x},$
@@
$P_{n}:=\left\{v\in V_{n}:\text{ for any }0\leqslant k\leqslant n,\leavevmode\nobreak\ \eta_{v,n}(k)<\frac{km_{n}}{n}\right\}.$
@@
#$\#\{v\in V_{n}:\eta_{v,n}(n)\geqslant m_{n}-x\}\geqslant\Delta_{n,x}.$ (16)
@@
$\mathbb{P}(H_{v,n}(x))$ $\gg n^{3/2}\rho^{-n}\mathbb{E}_{\mathbb{Q}}\left[e^{-\lambda_{n}(\eta_{v,n}(n)-m_{n})}{\mathbbm{1}}_{H_{v,n}(x)}\right]$
$\gg n^{3/2}\rho^{-n}\mathbb{E}_{\mathbb{Q}}\left[e^{c_{2}x}{\mathbbm{1}}_{H_{v,n}(x)}\right]=n^{3/2}\rho^{-n}e^{c_{2}x}\mathbb{Q}(H_{v,n}(x)).$
By the ballot theorem, in the form of Lemma 2.1 of [13], We have
$\mathbb{Q}(H_{v,n}(x))\gg n^{-3/2}\max(x-1,1)\gg n^{-3/2}x.$
@@
Next we estimate the second moment $\mathbb{E}[\Delta_{n,x}^{2}]$. Observe that by (A1) (see (29) of [13]),
@@
where $v\sim_{s}w$ means that the distance of $v$ and $w$ in the genealogical tree is equal to $2s$. To bound the latter probability we condition on the location of the common ancestor of $v$ and $w$, $\eta_{v,n}(n-s)$. We have $\text{for }v\sim_{s}w$,
@@
$\leqslant\sum_{j=0}^{\infty}\mathbb{P}\left(\eta_{v,n}(k)\leqslant\frac{km_{n}}{n}\text{ for }k\leqslant n-s,\leavevmode\nobreak\ \eta_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}\in[-(j+1),-j)\right)$
×$\hskip 113.81102pt\times\left(\sup_{y\in[-(j+1),-j)}P(n,s,x,y)\right)^{2},$
@@
$P(n,s,x,y)$ $:=\mathbb{P}\Bigg{(}\eta_{v,n}(n-s+\ell)\leqslant\frac{(n-s+\ell)m_{n}}{n}\text{ for }0\leqslant\ell\leqslant s,$
$\hskip 39.83368pt\eta_{v,n}(n)\in[m_{n}-(x+1),m_{n}-x)\mid\eta_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}=y\Bigg{)}.$
@@
$\hskip 14.22636pt\mathbb{P}\left(\eta_{v,n}(k)\leqslant\frac{km_{n}}{n}\text{ for }k\leqslant n-s,\leavevmode\nobreak\ \eta_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}\in[-(j+1),-j)\right)$
@@
×$\hskip 56.9055pt\times(\max(1,n-s))^{-3/2}\max\left((m_{n-s}-(n-s)m_{n}/n+j),1\right)$
@@
×$\hskip 56.9055pt\times\max\left(\big{(}j+\frac{(n-s)\log n}{n}-\log(n-s)\big{)},1\right)$

### v2

@@ 1.3 Proof strategy
In this subsection, we discuss the high-level intuition of the strategy we will follow, building from the case of $d=1$. Throughout, we condition on the survival event. We first assume that the jump distribution is spherically symmetric. For one-dimensional BRW, we expect the following parity relation between maximum and FPT. Recall that the asymptotic for the maximum $M_{n}$ at time $n$ is
@@
$\tau_{x}=t_{1}(x)+O_{\mathbb{P}}(1),\quad\text{ where }\quad t_{1}(x):=\frac{x}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log x,$ (9)
which we will prove in Proposition 7 below in Section 2.1. Recall (3). Let us decompose $A(x)=t_{1}(x)+t_{2}(x)$, where
$t_{2}(x):=A(x)-t_{1}(x)=\frac{d-1}{2c_{2}c_{1}}\,\log x.$ (10)
For a general dimension $d\geqslant 1$, we shall explain the extra term $t_{2}(x)$.
@@
A key step in these bounds is Proposition 8, which keeps track of the number of particles that are near the frontier (say, a distance of $\log n$ away) for a one-dimensional BRW. Roughly speaking, there are around $xe^{c_{2}x}$ particles that are of distance $x$ behind the frontier, where $x\ll\log n$. This amounts to generalizing the proof of the asymptotic of the maximum (8); see [bramson2016convergence]. A notable distinction is that unfortunately, one cannot conclude from the second moment method the lower bound for the number of particles near the frontier with overwhelming probability. Instead, one only concludes the lower bound with a uniformly positive probability. This gap can be filled using Theorem 2, which shows that a uniformly positive probability would already be sufficient.
A particularly nice feature of the spherical symmetry of the jumps is that the displacements in different directions are approximately independent (likewise, the same result Theorem 1 holds if we instead assume independent centered jumps in the $d$ directions). In particular, at time $n$, one roughly expects a proportion $O(n^{-(d-1)/2})$ of the particles that lie within unit distance from the origin in the last d1d-1 dimensions. In other words, one expects that the first passage event is realized around the time when $x^{(d-1)/2}$ particles reach a distance of $x$ for the $\mathbb{R}^{d}$-valued BRW projected onto the first dimension; see Figure 2 below. On the other hand, by (8), the extra time $t_{2}(x)$ from (10) pushes the frontier ahead by a distance of approximately $c_{1}t_{2}(x)=(d-1)(\log x)/(2c_{2})$, which amounts to having approximately $x^{(d-1)/2}$ extra particles in view of Proposition 8.
@@
If the jump distribution is not spherically symmetric, applying similar arguments above (with Proposition 8 replaced by Theorem B of [biggins1979growth] and Proposition 10 by Proposition 17; we omit the details here) will lead to
@@
where the corresponding upper and lower bounds for $\tau_{x}$ are separated by an order of $\log x$. To pin down the coefficient of the logarithmic correction term, we extend the arguments for the maximum of one-dimensional BRW ([bramson2016convergence]) with a modified barrier. The modified barrier event lies in $\mathbb{R}^{d}$ and hence results on random walks in cones ([denisov2015random]) are applicable. Note that the same proof also applies for the spherically symmetric case, but would yield less precise asymptotics compared to (4) (an error of $o_{\mathbb{P}}(\log x)$ instead of $O_{\mathbb{P}}(\log\log x)$).
1.4 Applications to polymer physics
Our interest in estimating the FPT of spatial branching processes stems from their relevance to polymer physics. The mechanical behavior of polymer networks appears to be controlled by the shortest paths between distant nodes in the polymer network, as shown by recent works [yin2020topological, yin2024network] using coarse-grained molecular dynamics (CGMD) simulations. The shortest path length (SP) serves as an indicator of how stretched the average load-bearing chains are in the polymer network. Theories based on SP go beyond traditional polymer models focused on elastic behaviors and aim to explain the inelastic behaviors exhibited by such materials.
The SP between (fixed) distant nodes is equivalent to the FPT of the BRW beginning from the origin (source node) and terminating at a desired point in space (destination node). Moreover, the occurrence of the cross-links along the polymer backbone mimics a Poisson process. As a result, a cross-link between two chains is equivalent to a branching event observed in the BRW genealogy tree. The agreement of the mean FPT from the BRW theory with the mean SP from the CGMD simulation results has been shown in our recent work [zhang2024modeling]. Furthermore, the numerical implementation of the BRW model shows an excellent agreement between its FPT distribution and the SP distribution from the CGMD simulation. Our earlier work initially focused on constant length jumps with the BRW model where there existed a correlation between the components of a jump step in the $d$ dimensions by virtue of being sampled from the unit sphere $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\left\lVert\mathbf{x}\right\rVert=1\}$. This was later shown to be represented equally well by the Gaussian BRW and BBM in which the components of each jump step are independent in the $d$ dimensions. However, the analysis was restricted to spherically symmetric jumps to accurately model the as-prepared polymer network prior to deformation.
@@ 2 Proofs of main results
This section is devoted to the proofs of the main results. We will first analyze the simple case of $d=1$ in Section 2.1, where a precise asymptotic of the FPT follows directly from inverting the maxima asymptotic (8). We then collect the necessary tools in Sections 2.2.1 and 2.2.2, and prove Theorems 1 and 2 in Sections 2.2.3 and 2.3 respectively. Most of the arguments before and including Section 2.2.1 closely follow the approach of [bramson2016convergence]. [^7] The results in [bramson2016convergence] are stated for the case $p_{0}=0$, but the general case follows in a similar way. The case of non-spherically symmetric jumps (Theorem 3) will be addressed in Section 2.4, where we also summarize preliminary results on random walks that are constrained to stay in a cone. Section 2.5 presents the proofs of Corollaries 4 and 5.
We begin by introducing some necessary notations. Following [bramson2016convergence], we denote by $V_{n}$ the offspring on the $n$-th generation. Let $\{\bm{\eta}_{v,n}(k)\}_{k=0,\dots,n}$ denote the $d$-dimensional random walk where $\bm{\eta}_{v,n}(k)$ is the position of the $k$-th generation for the path in the tree leading to a vertex $v\in V_{n}$. We also let $\eta^{(1)}_{v,n}(k)$ denote the first coordinate of $\bm{\eta}_{v,n}(k)$. When $d=1$, these notions coincide. For deterministic quantities or functions $A,B$, We use Vinogradov’s symbol $A\ll B$ (or $A=O(B)$) to denote $|A|\leqslant CB$ with some constant $C>0$ that depends only on the distribution of $\bm{\xi}$} and the dimension $d$. We write $A\asymp B$ if $A\ll B\ll A$. The notation $\left\lVert\cdot\right\rVert$ always denotes the Euclidean norm. We use $B_{\mathbf{x}}(r)$ to denote the shifted ball of radius $r$ centered at $\mathbf{x}\in\mathbb{R}^{d}$, $B_{\mathbf{x}}=B_{\mathbf{x}}(1)$, and $B_{x}=B_{(x,\mathbf{0})}$ for $x\in\mathbb{R}$. From here on, without loss of generality, we will assume that quantities that tend to infinity are integers. The arguments can be made rigorous by writing floor functions everywhere, but we have chosen not to do so for the simplicity of the notation.
@@ 2.1 Warming up in dimension one
For $\beta\geqslant 1$ and $n\in\mathbb{N}$ we define the event
$G_{n,\beta}:=\bigcup_{v\in V_{n}}\bigcup_{0\leqslant k\leqslant n}\left\{\eta^{(1)}_{v,n}(k)\geqslant\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\right\}.$
Lemma 6 (Lemma 2.4 of [bramson2016convergence]).
Assume conditions (A1), (A2), and (A4), and assume that the jump distribution is non-lattice. It holds that for all $n\in\mathbb{N}$, $\mathbb{P}(G_{n,\beta})\ll\beta e^{-c_{2}\beta}$ for $\beta\geqslant 1$.
@@
Suppose that a one-dimensional BRW satisfies conditions (A1), (A2), and (A4), $p_{0}=0$ (no extinction) and that the jump distribution is non-lattice. We have
$\tau_{x}=t_{1}(x)+O_{\mathbb{P}}(1)=\frac{x}{c_{1}}+\frac{3}{2c_{2}c_{1}}\,\log x+O_{\mathbb{P}}(1).$
@@
For the lower bound, we fix $\varepsilon>0$ and pick $\beta$ large enough such that $\mathbb{P}(G_{n,\beta})<\varepsilon$ for all $n\in\mathbb{N}$, by means of Lemma 6. In other words, with a probability larger than $1-\varepsilon$, the whole BRW stays below the curve
@@
$\max_{1\leqslant k\leqslant t_{1}(x)-C_{1}}\Big(\frac{km_{t_{1}(x)-C_{1}}}{t_{1}(x)-C_{1}}+\beta+\frac{4}{c_{2}}(\log(\min(k,t_{1}(x)-C_{1}-k)))_{+}\Big)\leqslant x.$
This yields $\mathbb{P}(\tau_{x}\leqslant t_{1}(x)-C_{1})\leqslant\mathbb{P}(G_{n,\beta})<\varepsilon$, proving the lower bound.
For the upper bound, we fix $\varepsilon>0$. Pick $C_{2}>0$ such that $\mathbb{P}(M_{t_{1}(x)+C_{2}}>x)>1-\varepsilon/3$ uniformly in $x$ by (8) and the convergence in law of the recentered maximum (Theorem 1.1 of [bramson2016convergence]). Using the same analysis for the lower bound, there exists $C_{3}>0$ such that $\mathbb{P}(M_{t_{1}(x)+C_{2}}>x+C_{3})<\varepsilon/3$. Therefore, with probability $1-2\varepsilon/3$, there is $v\in V_{t_{1}(x)+C_{2}}$ such that $x\leqslant\eta^{(1)}_{v,t_{1}(x)+C_{2}}(t_{1}(x)+C_{2})\leqslant x+C_{3}$. Since the distribution of $\xi$} is non-lattice, by the Chung–Fuchs theorem (Theorem 5.4.8 of [durrett2019probability]), the random walk with step distribution $\xi$ initiated from $v$ is recurrent, and hence there exists $C_{4}>0$ such that the descendants of $v$ reaches [x1/2,x+1/2][x-1/2,x+1/2] with probability $1-\varepsilon/3$ within time $C_{4}$. Combining the above leads to $\mathbb{P}(\tau_{x}>t_{1}(x)+C_{2}+C_{4})<\varepsilon$, as desired. ∎
@@ 2.2 Proof of Theorem 1
We now present the proof of Theorem 1 for $d\geqslant 2$, provided that Theorem 2 holds true. The proof of Theorem 2 will be the focus of the next section. The following auxiliary results are necessary as explained in our proof strategy.
@@ 2.2.1 The number of particles near maximum
The goal of this section is to prove the following result on the number of particles near the frontier of a one-dimensional BRW, which might be of independent interest.
@@
Assume (A1)–(A4). There exists $L>1$ depending only on the law of the BRW such that the following holds conditioned upon survival. Given any $\varepsilon>0$, there exists $C>0$ independent from $n$ and $x$ such that uniformly for $n$ large enough and for $x\in[2,\sqrt{n}]$,
$\mathbb{P}\left(\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}>Cxe^{c_{2}x}\mid S\right)<\varepsilon$ (12)
@@
$\mathbb{P}\left(\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}>\frac{1}{C}xe^{c_{2}x}\mid S\right)>\frac{1}{L}.$ (13)
@@
Proposition 8 may be compared against [biggins1979growth, rouault1993precise], where it is proven that for a location that is $\Omega(n)$ away from $m_{n}$, the number of particles therein has the same order as its expectation. Proposition 8 shows that this is not the case for locations that are $O(\sqrt{n})$ near $m_{n}$. On the other hand, the number of particles that are $O(\sqrt{n})$ away from the origin can be described by the BRW central limit theorem [biggins1990central]. An interesting question would be to analyze the phase transitions between those regimes. We also expect that (13) holds if we replace its right-hand side by $1-\varepsilon$, but the current version suffices for our purpose.
In the rest of this section, we assume that $\mathbb{P}$ is the probability measure conditioned upon the survival event $S$. Following the setup of Section 2.2 of [bramson2016convergence], we introduce $\lambda_{n}$ which is the value of $\lambda$ where the supremum of (1) is attained with $x=m_{n}/n$. By (15) of [bramson2016convergence], it holds $0\leqslant c_{2}-\lambda_{n}\ll(\log n)/n$. For a given $v\in V_{n}$, let $\mathbb{Q}^{(n)}$ be a probability measure on length-$n$ random walk paths defined by
$\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}^{(n)}}:=e^{-\lambda_{n}(\eta^{(1)}_{v,n}(n)-m_{n})-nI(m_{n}/n)}\asymp n^{3/2}\rho^{-n}e^{-\lambda_{n}(\eta^{(1)}_{v,n}(n)-m_{n})}.$ (14)
It follows that under $\mathbb{Q}^{(n)}$, $\{\eta^{(1)}_{v,n}(k)-km_{n}/n\}_{k=0,\dots,n}$ is a mean zero random walk. Let
$Q_{n,\beta}:=\left\{v\in V_{n}:\text{ for any }0\leqslant k\leqslant n,\penalty 10000\ \eta^{(1)}_{v,n}(k)<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\right\}$
@@
$g(n,\beta,x):={\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}]}.$ (15)
@@
$\mathbb{E}\left[\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}{\mathbbm{1}}_{G_{n,\beta}^{c}}\right]\ll\beta(x+\beta)e^{c_{2}x}.$
@@
$\mathbb{P}\left(G_{n,\beta}^{c}\cap\left\{\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}>L\beta(x+\beta)e^{c_{2}x}\right\}\right)<\frac{\varepsilon}{2}.$
@@
Let $\{S_{k}\}_{1\leqslant k\leqslant n}$ be a random walk with increment $\xi^{(1)}$. By the many-to-one formula, we have
$\hskip 14.22636pt\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}]$
$=\rho^{n}\mathbb{P}\Big(S_{k}<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\text{ for }1\leqslant k\leqslant n;\,S_{n}\geqslant m_{n}-x\Big)$
$=\sum_{j=0}^{x+\beta-1}\rho^{n}\mathbb{P}\Big(S_{k}<\frac{km_{n}}{n}+\beta+\frac{4}{c_{2}}(\log(\min(k,n-k)))_{+}\text{ for }1\leqslant k\leqslant n;$
$\hskip 113.81102ptS_{n}\in[m_{n}+\beta-j-1,m_{n}+\beta-j)\Big).$
By substituting $i=\beta-j$, the latter probability is precisely $\chi_{n,n}^{\mathbb{P}}(i)$ defined below (15) in [bramson2016convergence]. It is proved in (18) therein the estimate
@@
$\mathbb{E}[\#\{v\in Q_{n,\beta}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}]\ll\sum_{j=0}^{x+\beta}\beta(j+2)e^{-c_{2}(\beta-j)}\ll\beta(x+\beta)e^{c_{2}x},$
@@
$P_{n}:=\left\{v\in V_{n}:\text{ for any }0\leqslant k\leqslant n,\penalty 10000\ \eta^{(1)}_{v,n}(k)<\frac{km_{n}}{n}\right\}.$
@@
$\#\{v\in V_{n}:\eta^{(1)}_{v,n}(n)\geqslant m_{n}-x\}\geqslant\Delta_{n,x}.$ (16)
@@
$\mathbb{P}(H_{v,n}(x))$ $\gg n^{3/2}\rho^{-n}\mathbb{E}_{\mathbb{Q}^{(n)}}\left[e^{-\lambda_{n}(\eta^{(1)}_{v,n}(n)-m_{n})}{\mathbbm{1}}_{H_{v,n}(x)}\right]$
$\gg n^{3/2}\rho^{-n}\mathbb{E}_{\mathbb{Q}^{(n)}}\left[e^{c_{2}x}{\mathbbm{1}}_{H_{v,n}(x)}\right]=n^{3/2}\rho^{-n}e^{c_{2}x}\mathbb{Q}^{(n)}(H_{v,n}(x)).$
By the ballot theorem, in the form of Lemma 2.1 of [bramson2016convergence], We have
$\mathbb{Q}^{(n)}(H_{v,n}(x))\gg n^{-3/2}\max(x-1,1)\gg n^{-3/2}x.$
@@
Next we estimate the second moment $\mathbb{E}[\Delta_{n,x}^{2}]$. Observe that by (A1) (see (29) of [bramson2016convergence]),
@@
where $v\sim_{s}w$ means that the distance of $v$ and $w$ in the genealogical tree is equal to $2s$. To bound the latter probability we condition on the location of the common ancestor of $v$ and $w$, $\eta^{(1)}_{v,n}(n-s)$. We have $\text{for }v\sim_{s}w$,
@@
$\leqslant\sum_{j=0}^{\infty}\mathbb{P}\left(\eta^{(1)}_{v,n}(k)\leqslant\frac{km_{n}}{n}\text{ for }k\leqslant n-s,\penalty 10000\ \eta^{(1)}_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}\in[-(j+1),-j)\right)$
$\hskip 113.81102pt\times\left(\sup_{y\in[-(j+1),-j)}P(n,s,x,y)\right)^{2},$
@@
$P(n,s,x,y)$ $:=\mathbb{P}\Bigg(\eta^{(1)}_{v,n}(n-s+\ell)\leqslant\frac{(n-s+\ell)m_{n}}{n}\text{ for }0\leqslant\ell\leqslant s,$
$\hskip 34.14322pt\eta^{(1)}_{v,n}(n)\in[m_{n}-(x+1),m_{n}-x)\mid\eta^{(1)}_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}=y\Bigg).$
@@
$\hskip 14.22636pt\mathbb{P}\left(\eta^{(1)}_{v,n}(k)\leqslant\frac{km_{n}}{n}\text{ for }k\leqslant n-s,\penalty 10000\ \eta^{(1)}_{v,n}(n-s)-\frac{(n-s)m_{n}}{n}\in[-(j+1),-j)\right)$
@@
$\hskip 56.9055pt\times(\max(1,n-s))^{-3/2}\max\left((m_{n-s}-(n-s)m_{n}/n+j),1\right)$
@@
$\hskip 56.9055pt\times\max\left(\big(j+\frac{(n-s)\log n}{n}-\log(n-s)\big),1\right)$